OpenStudy (anonymous):

Write the given expression as an algebraic expression in x. cos(2 cos−1 x)

OpenStudy (freckles):

you will need the double angle identity

OpenStudy (freckles):

\[\cos(2a)=\cos^2(a)-\sin^2(a)\] let a=arccos(x) then cos(a)=x draw a right triangle using the equation cos(a)=x Then find any missing measurements by pythagorean theorem

OpenStudy (anonymous):

I am unsure how to make the triangle